# Erasing Without Remembering: Safeguarding Knowledge Forgetting in Large Language Models

### v1

@@ 7.3 Tuning Coefficient $C$
We investigate the impact of the Tuning Coefficient $C$, which varies from $0.0$ to $1.0$ in increments of $0.1$. The experiments are conducted on the TOFU Forget01 dataset using Phi-1.3B, with all other parameters held constant. As shown in Tab. XII, the best model utility is achieved when $C=0.0$, where only the corrupted-run probability distribution is used. While this setting maintains high model utility, the unlearning effect is insufficient, as the top-ranked token in the clean-run probability distribution is not fully suppressed. As $C$ increases, the unlearning effect improves, reaching its peak when $C=0.4$. Then it begins to fluctuate as $C$ continues to increase. Correspondingly, model utility decreases with larger values of $C$, which is expected, as higher $C$ values subtract more information from $p(y|x^{\prime})$, potentially disrupting the distribution of irrelevant knowledge. To balance model utility with effective unlearning, we select $C=0.1$ for our experiments.
TABLE XIII: A comparison of perturbing a single subject in sentences (referred to as $\textsc{PerMU}_{single}$) versus multiple subjects on the Harry Potter dataset. The results indicate that perturbing just one subject achieves competitive performance compared to perturbing all subjects, especially for smaller models.
Model Dataset Forget data Retain data Real Authors Real World MU↑ FRT↑
Metric RG↓ Pr↓ F1↓ RG↑ Pr↑ F1↑ RG↑ Pr↑ TR↑ RG↑ Pr↑ TR↑
Phi (1.3B) PerMU 65.21 35.75 63.02 82.58 66.08 81.61 60.90 48.89 59.98 67.09 52.47 64.91 63.20 1.16
$\textsc{PerMU}_{single}$ 63.35 36.49 61.45 81.53 65.20 80.47 62.90 48.82 59.89 67.09 52.44 64.99 63.17 1.17
LLaMA (7B) PerMU 71.88 56.36 69.54 92.42 88.88 91.54 83.25 69.40 83.66 85.47 71.25 84.84 82.65 1.25
$\textsc{PerMU}_{single}$ 75.30 61.77 72.41 90.55 85.73 88.91 85.42 65.58 81.01 85.04 69.40 83.83 80.82 1.16
7.4 Multiple Subjects Scenarios
The simplest unlearning scenario involves a sentence built around only one subject of knowledge, as seen in datasets like TOFU and ZsRE. However, in real-world scenarios, sentences often contain multiple names or entities serving as subjects. This complexity poses greater challenges for the model to forget specific knowledge, as the information stored in language models is highly entangled. In this experiment, we investigate this more complex scenario and use the Harry Potter dataset, as sentences in novels often contain multiple subjects. For example, consider the sentence: “How many points were taken from Gryffindor due to Harry, Hermione, and Neville being caught out of bed?”. To ensure a fair comparison, we randomly perturb only one subject in the sentence, a method we denote as $\textsc{PerMU}_{single}$. In contrast, PerMU represents our original proposed approach, where all subject tokens in the sentence are perturbed. The experiments are conducted on both Phi-1.3B and LLaMA2-7B, with all other parameters kept constant. As shown in Tab. XIII, to our surprise, $\textsc{PerMU}_{single}$ achieves competitive unlearning performance while maintaining considerable model utility. This phenomenon is particularly noticeable when using Phi-1.3B, where $\textsc{PerMU}_{single}$ even surpasses PerMU relatively by up to 2.9% in Forget ROUGE. Although $\textsc{PerMU}_{single}$ may become slightly less effective than PerMU with model scaling, it still outperforms other baselines presented in Tab.VI. In summary, perturbing all subject tokens with PerMU demonstrates more robust unlearning ability, particularly in larger models. However, for smaller models, perturbing only one subject in the sentence with $\textsc{PerMU}_{single}$ can still ensure promising unlearning performance while preserving model utility.
@@ 7.5 Discrete-Token Level Perturbation
Apart from adding random noise to the subject token embeddings, PerMU can also be implemented by perturbing the subject words at the discrete-token level, denoted as $\textsc{PerMU}_{dis}$. In this experiment, we evaluate the unlearning performance of $\textsc{PerMU}_{dis}$ on the TOFU dataset while keeping all other parameters constant. The perturbation type is randomly chosen from deleting, altering, or adding letters to the subject words. As shown in Tab. XIV, $\textsc{PerMU}_{dis}$ exhibits exceptional unlearning capability, outperforming PerMU in Forget ROUGE across all datasets and models. Moreover, $\textsc{PerMU}_{dis}$ achieves an absolute 19.26% Forget ROUGE and 4.64% Forget Probability on Forget05 when using LLaMA2-7B, surpassing PerMU by up to 11.16% (30.42$\rightarrow$19.26) and 11.65% (16.29$\rightarrow$4.64), respectively, though with a slightly lower model utility of 0.92%. However, when tested on Phi-1.3B, $\textsc{PerMU}_{dis}$ does not consistently exhibit superior unlearning performance, and the model utility drops by up to 4.44% (50.94$\rightarrow$46.50) as the number of unlearning samples increases. Despite this, the FRT ratio of $\textsc{PerMU}_{dis}$ still outperforms other baselines presented in Tab. II. In summary, perturbing subject words at the discrete-token level can also prevent the model from recalling the fact and generate fact-unrelated probability distributions, thus achieving unlearning. Both embedding-layer and discrete-token-level noise methods can achieve effective unlearning but result in different trade-offs. Given that the knowledge retention ability of $\textsc{PerMU}_{dis}$ may decline as the amount of forgotten data increases, we choose to add noise at the embedding layer as a more promising alternative.
TABLE XIV: Experimental results for implementing perturbation at the discrete-token level to subject words, denoted as $\textsc{PerMU}_{dis}$. Surprisingly, $\textsc{PerMU}_{dis}$ demonstrates a stronger unlearning effect on some datasets, particularly achieving 19.26% in Forget Rouge and 4.64% in Probability on the Forget01 dataset using LLaMA2-7B. These results highlight the adaptability of our approach across different dimensions.
@@
Phi (1.3B) Forget01 PerMU 46.26 27.11 64.65 67.74 79.50 44.96 41.90 37.85 44.04 76.14 41.75 50.39 50.14 1.37
$\textsc{PerMU}_{dis}$ 35.46 27.53 53.67 84.31 89.36 48.18 46.23 38.26 46.71 76.10 40.91 49.67 52.72 1.67
@@
$\textsc{PerMU}_{dis}$ 36.94 36.76 60.86 42.37 58.66 41.05 45.82 37.43 44.93 78.40 40.73 48.92 46.50 1.26
Forget10 PerMU 46.55 41.93 57.51 81.59 86.17 47.10 35.23 37.64 45.17 75.28 41.25 49.92 50.07 1.13
$\textsc{PerMU}_{dis}$ 44.90 57.06 55.92 46.25 63.60 43.77 49.35 36.89 44.34 73.65 39.78 47.41 47.39 0.93
LLaMA (7B) Forget01 PerMU 30.42 16.29 74.11 86.67 88.55 43.51 92.80 51.82 66.34 89.17 49.28 63.02 65.06 2.79
$\textsc{PerMU}_{dis}$ 19.26 4.64 73.92 83.72 89.11 42.39 91.00 51.19 66.05 87.46 48.78 62.66 64.14 5.37
Forget05 PerMU 33.66 39.23 64.39 83.63 88.24 41.89 91.30 52.57 68.21 89.60 49.77 63.94 64.89 1.78
$\textsc{PerMU}_{dis}$ 30.39 36.55 65.45 72.08 79.45 41.19 91.50 52.61 68.67 88.32 50.15 64.52 63.38 1.89
Forget10 PerMU 44.60 47.58 67.25 94.14 94.68 41.23 90.30 51.66 66.56 88.75 48.48 62.10 64.80 1.41
$\textsc{PerMU}_{dis}$ 40.32 48.68 62.45 70.72 79.47 41.56 91.50 53.81 70.03 87.89 50.86 65.92 63.93 1.44
TABLE XV: A comparison of different retain loss settings. w/o GDR refers to using the vanilla forget loss, while w/ KLR indicates the combination with KLR. Overall, PerMU effectively preserves the model’s utility.
@@ 7.6 Different Retain Loss
We investigate the impact of different retain loss functions while keeping other parameters fixed. The experiments are conducted on the TOFU Forget10 dataset using LLaMA2-7B. PerMU typically employs the vanilla forget loss combined with GDR, with a retain weight of $RW=1$. Here, w/o GDR refers to using the vanilla forget loss without any retain loss, while w/ KLR denotes the combination of the vanilla forget loss with KLR, applying the same retain weight. As shown in Tab. XV, using the vanilla forget loss achieves strong unlearning performance but may slightly impair model utility. While incorporating KLR can improve model utility, the enhancement is less significant compared to using GDR as the retain loss. Therefore, we primarily adopt GDR as the retain loss in our experiments.
@@ 8 Conclusions
In this paper, we strive to safeguard LLM-based model unlearning by enhancing generalisation capabilities, ensuring that unlearned models do not retain any memory of the targeted knowledge. To this end, we introduce UGBench, a comprehensive benchmark for assessing the generalisation performance of unlearning methods, and propose PerMU, a novel approach that addresses two key issues from UGBench. Our findings underscore the importance of moving beyond superficial forgetting, paving the way for safer and more thorough LLM-based unlearning. In our future work, we plan to incorporate human evaluations in assessments.

### v2

@@ 7.7 Tuning Coefficient $C$
We investigate the impact of the Tuning Coefficient $C$, which varies from $0.0$ to $1.0$ in increments of $0.1$. The experiments are conducted on the TOFU Forget01 dataset using Phi-1.3B, with all other parameters held constant. As shown in Tab. XVI, the best model utility is achieved when $C=0.0$, where only the corrupted-run probability distribution is used. While this setting maintains high model utility, the unlearning effect is insufficient, as the top-ranked token in the clean-run probability distribution is not fully suppressed. As $C$ increases, the unlearning effect improves, reaching its peak when $C=0.4$. Then it begins to fluctuate as $C$ continues to increase. Correspondingly, model utility decreases with larger values of $C$, which is expected, as higher $C$ values subtract more information from $p(y|x^{\prime})$, potentially disrupting the distribution of irrelevant knowledge. To balance model utility with effective unlearning, we select $C=0.1$ for our experiments.
@@ 7.8 Discrete-Token Level Perturbation
Apart from adding random noise to the subject token embeddings, PerMU can also be implemented by perturbing the subject words at the discrete-token level, denoted as $\textsc{PerMU}^{\dagger}_{dis}$. In this experiment, we evaluate the unlearning performance of $\textsc{PerMU}^{\dagger}_{dis}$ on the TOFU dataset while keeping all other parameters constant. The perturbation type is randomly chosen from deleting, altering, or adding letters to the subject words.
As shown in Tab. XVII, $\textsc{PerMU}^{\dagger}_{dis}$ exhibits exceptional unlearning capability, outperforming $\textsc{PerMU}^{\dagger}$ in Forget ROUGE across all datasets and models. Moreover, $\textsc{PerMU}^{\dagger}_{dis}$ achieves an absolute 19.26% Forget ROUGE and 4.64% Forget Probability on Forget05 when using LLaMA2-7B, surpassing $\textsc{PerMU}^{\dagger}$ by up to 11.16% (30.42$\rightarrow$19.26) and 11.65% (16.29$\rightarrow$4.64), respectively, though with a slightly lower model utility of 0.92%. However, when tested on Phi-1.3B, $\textsc{PerMU}^{\dagger}_{dis}$ does not consistently exhibit superior unlearning performance, and the model utility drops by up to 4.44% (50.94$\rightarrow$46.50) as the number of unlearning samples increases. Despite this, the FRT ratio of $\textsc{PerMU}^{\dagger}_{dis}$ still outperforms other baselines. In summary, perturbing subject words at the discrete-token level can also prevent the model from recalling the fact and generate fact-unrelated probability distributions, thus achieving unlearning. Both embedding-layer and discrete-token-level noise methods can achieve effective unlearning but result in different trade-offs. Given that the knowledge retention ability of $\textsc{PerMU}^{\dagger}_{dis}$ may decline as the amount of forgotten data increases, we choose to add noise at the embedding layer as a more promising alternative.
TABLE XVII: Experimental results for implementing perturbation at the discrete-token level to subject words, denoted as $\textsc{PerMU}_{dis}$. Surprisingly, $\textsc{PerMU}^{\dagger}_{dis}$ demonstrates a stronger unlearning effect on some datasets, particularly achieving 19.26% in Forget Rouge and 4.64% in Probability on the Forget01 dataset using LLaMA2-7B. These results highlight the adaptability of our approach across different dimensions.
@@
Phi (1.3B) Forget01 $\textsc{PerMU}^{\dagger}$ 46.26 27.11 64.65 67.74 79.50 44.96 41.90 37.85 44.04 76.14 41.75 50.39 50.14 1.37
$\textsc{PerMU}^{\dagger}_{dis}$ 35.46 27.53 53.67 84.31 89.36 48.18 46.23 38.26 46.71 76.10 40.91 49.67 52.72 1.67
@@
$\textsc{PerMU}^{\dagger}_{dis}$ 36.94 36.76 60.86 42.37 58.66 41.05 45.82 37.43 44.93 78.40 40.73 48.92 46.50 1.26
Forget10 $\textsc{PerMU}^{\dagger}$ 46.55 41.93 57.51 81.59 86.17 47.10 35.23 37.64 45.17 75.28 41.25 49.92 50.07 1.13
$\textsc{PerMU}^{\dagger}_{dis}$ 44.90 57.06 55.92 46.25 63.60 43.77 49.35 36.89 44.34 73.65 39.78 47.41 47.39 0.93
LLaMA (7B) Forget01 $\textsc{PerMU}^{\dagger}$ 30.42 16.29 74.11 86.67 88.55 43.51 92.80 51.82 66.34 89.17 49.28 63.02 65.06 2.79
$\textsc{PerMU}^{\dagger}_{dis}$ 19.26 4.64 73.92 83.72 89.11 42.39 91.00 51.19 66.05 87.46 48.78 62.66 64.14 5.37
Forget05 $\textsc{PerMU}^{\dagger}$ 33.66 39.23 64.39 83.63 88.24 41.89 91.30 52.57 68.21 89.60 49.77 63.94 64.89 1.78
$\textsc{PerMU}^{\dagger}_{dis}$ 30.39 36.55 65.45 72.08 79.45 41.19 91.50 52.61 68.67 88.32 50.15 64.52 63.38 1.89
Forget10 $\textsc{PerMU}^{\dagger}$ 44.60 47.58 67.25 94.14 94.68 41.23 90.30 51.66 66.56 88.75 48.48 62.10 64.80 1.41
$\textsc{PerMU}^{\dagger}_{dis}$ 40.32 48.68 62.45 70.72 79.47 41.56 91.50 53.81 70.03 87.89 50.86 65.92 63.93 1.44
TABLE XVIII: A comparison of different retain loss settings. w/o GDR refers to using the vanilla forget loss, while w/ KLR indicates the combination with KLR. Overall, $\textsc{PerMU}^{\dagger}$ effectively preserves the model’s utility.
@@ 7.9 Different Retain Loss
We investigate the impact of different retain loss functions while keeping other parameters fixed. The experiments are conducted on the TOFU Forget10 dataset using LLaMA2-7B. $\textsc{PerMU}^{\dagger}$ typically employs the forget loss combined with GDR, with a retain weight of $RW=1$. Here, w/o GDR refers to using the vanilla forget loss without any retain loss, while w/ KLR denotes the combination of the vanilla forget loss with KLR, applying the same retain weight. As shown in Tab. XVIII, using the vanilla forget loss achieves strong unlearning performance but may slightly impair model utility. While incorporating KLR can improve model utility, the enhancement is less significant compared to using GDR as the retain loss. Therefore, we primarily adopt GDR as the retain loss in our experiments.
@@ 8 Conclusions
In this paper, we advance LLM-based knowledge forgetting by shifting the focus from superficial forgetting to the more challenging task of implicit knowledge forgetting. To this end, we introduce UGBench, a comprehensive benchmark for evaluating implicit unlearning, and propose PerMU, a novel approach that addresses two core challenges identified by UGBench. Our results highlight the need for more generalised and thorough unlearning in LLMs. While automatic evaluation metrics offer reliable assessments of model performance, we plan to incorporate human evaluations in future work to further enhance the evaluation process. Overall, we believe our contributions offer meaningful insights into effective knowledge unlearning in NLP, with broader implications for related areas such as model editing, knowledge conflict resolution, reasoning, and other diverse fields.